How would you determine the critical volume of a gas?
1
Expert's answer
2019-10-24T08:00:19-0400

"T_c=\frac{8a}{27Rb}"

"P_c=\frac{a}{27b^2}"

Hence, using "P_c\times V_c=R\times T_c"

Critical volume "V_c=3b"